OpenStudy (anonymous):

f(x)= .5(x-1)^3. find f'(0) PLEASE SHOW STEPS!

OpenStudy (anonymous):

15

OpenStudy (hoblos):

f'(x)= 15(x-1)^2 f'(0)= 15(-1)^2 = 15